Solve the following system of equations.
1) 3x+5y=7
2) 2x-3y=1

Answer by richwmiller(17219)

3x+5y=7
2x-3y=1
multiply the first by 2 and the second by 3
6x+10y=14
6x-9y=3
subtract them
19y=11
y=11/19
2x-3(11/19)=1
2x-33/19=19/19
add 33/19 to both sides
2x=52/19
x=26/19
check
3(26/19)+5(11/19)=7
78+55=7*19
133=133
ok